Given utility function: U(x, y) = X1/3Y2/3, Px = 2, Py = 5 and M = 400, find:

a. The demand equation for X and Y.

b. The utility maximizing levels of X and Y.

c. The maximum utility.

d. The MRSx, y at the optimum level.


1
Expert's answer
2022-01-27T11:00:46-0500

U="(X^ \frac {1}{3}Y^\frac{2}{3})"

I= 400

"_x= 2"

"P_y= 5"

a)Demand functions

For utility maxmization


"\frac{\frac{1}{3}X^\frac{-2}{3}Y^\frac{2}{3}}{\frac{2}{3}X^\frac{1}{3}Y^\frac{-1}{3}}=\frac {P_x}{P_y}"


"\frac{\frac {1}{3}Y}{\frac{2}{3}X}=\frac {P_x}{P_y}"


"\frac{2}{3}P_xX=\frac{1}{3}P_yY"


X= "\frac{P_yY}{2P_x}" 


"2P_xX=P_yY"


"Y=\frac{2P_xX}{P_y}" 

Plug into the budget constraint


m= "P_xX+P_yY"

m="P_x(\frac{P_yY}{2P_x})+P_yY"

2m="3P_yY"


"Y^*= \frac{2m}{3P_y}"

m="P_xX+P_yY"

"m=P_xX+P_Y(\frac{2P_xX}{P_y})"

m"=P_xX+2P_xX"

X*="\frac{m}{3P_x}"


b) Utility Maximizing levels

"Y^*= \frac{2\times 400}{3\times 5}=\frac{800}{15}= 53.33"


c) Maximum utility of consuming the two goods

= 66.3=120

d) MRS

"MRS_(xy)= \frac{Mu_x}{Mu_y}= \frac{\frac{1}{3}X^\frac{-2}{3}Y^\frac{2}{3}}{\frac{2}{3}X^\frac{1}{3}Y^\frac{-1}{3}}"

"=\frac{\frac{1}{3}Y}{\frac{2}{3}X}=\frac{Y}{2X}="

"=\frac{53.33}{66.67\times 2}=\frac{53.33}{134.34}=0.4"